Evaluate the integral of cot (√x+1)/(√x+1)dx


1
Expert's answer
2021-03-07T17:33:31-0500

Solution.


"\int{\frac{\cot{\sqrt{x+1}}}{\sqrt{x+1}}}dx=
\newline
=2\ln{\sin{\sqrt{x+1}}}+C."